OpenStudy (jazzyluvsu123):

1/5x=3/10 i do Divide 3/10 / 1/5 ?? do i flip the 1/5 to 5/1 ??

OpenStudy (jazzyluvsu123):

@mathmaticallychallenged

OpenStudy (anonymous):

1/5x = 3/10 x = 5/1(3/10) x = 15/10

OpenStudy (jazzyluvsu123):

Simplified my answer would be 3/2 ?? @azirkelb

OpenStudy (anonymous):

yes